OpenStudy (anonymous):

simplify the expression ln(e^3/x^4)

OpenStudy (nenadmatematika):

ln(e^3)-ln(x^4)=3-4ln(x)

OpenStudy (nenadmatematika):

I used the rule ln(A^B)=B*ln(A) and ln(A/B)=lnA-lnB